# One-Shot Imitation under Mismatched Execution

### v1

@@ 3            $\pi_{\theta}\leftarrow\texttt{Update-Policy}(a_{t},\pi_{\theta}(s_{t},\mathbf{z_{R}}))$
Algorithm 2 RHyME: Retrieval for Hybrid imitation under Mismatched Execution
IV-B RHyME: Retrieval for Hybrid imitation under Mismatched Execution
Training Overview. Algorithm 2 provides an overview of our hybrid learning approach to train a robot policy that learns to condition on mismatched demonstration videos. The policy is trained on robot data that has robot trajectory $\mathbf{\xi_{R}}$ and the robot video $\mathbf{v_{R}}$. The policy is trained in a hybrid fashion over two modes: (1) on robot video and 2) on imagined demonstration video. For mode (1), the robot video is encoded $\mathbf{z_{R}}=\phi(\mathbf{v_{R}})$ using the learned video encoder $\phi$ that we discuss in Section. IV-A. The policy is then trained on robot video $\pi(a_{t}|s_{t},\mathbf{z_{R}})$. While the aligned embeddings enable a policy to transfer from robot video to demonstrator video, significant execution mismatches can result in a distribution mismatch between $\mathbf{z_{R}}$ and $\mathbf{z_{H}}$.
Hence we also train using mode (2), where we imagine a sequence of demonstrator embeddings $\mathbf{\hat{z}_{H}}$ that encodes how the true demonstrator carries out the same task specification as in the robot’s video $\mathbf{v_{R}}$. Using this imagined correspondence between robot actions and a demonstrator video, we can effectively learn a hybrid policy to take input embeddings generated from either a robot or demonstrator video.
@@ IV-B RHyME: Retrieval for Hybrid imitation under Mismatched Execution
Figure 3: Cross-Embodiment Vision Embeddings. (Left) Visualizing task embeddings. We use t-SNE to visualize cross-embodiment latent embeddings generated by robot and demonstrator when executing different tasks on all three datasets. (Right) TCC Failure Example: The robot and video clip are equivalent, but specific frames have high TCC losses. For example, a frame showing the robot performing the ’kettle’ action has a high loss due to its nearest neighbor in the video performing both ’kettle’ and ’light’ actions. This frame cycles back to the robot performing ’light’, which is mismatched.
@@
Imagining Demonstration Videos. Algorithm 1 details the retrieval process for imaging a sequence of demonstration embeddings. We first divide the robot’s long-horizon video embeddings into short-horizon windows that can be compared with the demonstrator’s play dataset. Then, we compare the robot’s short-horizon video embeddings with those generated from the play dataset, retrieving the one with the lowest sequence-level distance. Finally, we concatenate the retrieved video embeddings into one sequence of imagined long-horizon demonstration video embeddings. Integral to this retrieval process is the definition of a distance function $d(\mathbf{z_{R}},\mathbf{z_{H}})$ that can handle inputs of video sequences with potentially different lengths. We construct two methods to compute the distance.
@@
Method 1: Computing Distance via Optimal Transport. In this formulation, we define the distance we attempt to find matching between the sequence between robot and demonstrator video embeddings. We do this by calculating the total Wasserstein distance between the videos, i.e., the cost of the optimal transport plan that transfers one sequence of video embeddings into another. The robot’s embedding distribution is defined as $\rho_{R}=\{1/T,1/T,\dots,1/T\}$, and the demonstrator’s embedding distribution is defined as $\rho_{H}=\{1/T^{\prime},1/T^{\prime},\dots,1/T^{\prime}\}$, where $T$ and $T^{\prime}$ are the lengths of the video sequences respectively. The cost function for the transport is $C\in\mathbb{R}^{T\times T^{\prime}}$ where $C^{ij}$ is the cosine distance between the robot embedding $z_{R}^{i}$ and the demonstration embedding $z_{H}^{j}$. Our goal is to find the optimal assignment $M\in\mathbb{R}^{T\times T^{\prime}}$ that transports the distribution from $\rho_{R}$ to $\rho_{H}$ while minimizing the cost of the plan. Formally, we need to find $M^{*}=\operatorname*{arg\,min}\limits_{M^{*}}\sum_{i}\sum_{j}C^{ij}M^{ij}$. After solving the optimal transport assignment, the distance function is the cost of the plan, i.e., $d(\mathbf{z_{R}},\mathbf{z_{H}})=\sum_{i}\sum_{j}C^{ij}M^{*ij}$. In practice, we optimize an entropy-regularized version of this problem to find an approximate solution efficiently using the Sinkhorn-Knopp algorithm [39]. More details are provided in the appendix.
@@
Method 2: Computing Distance via Temporal-Cyclic Consistency (TCC). Originally proposed by [40], the TCC loss function can be used to match videos of a robot and its demonstrator [31, 24]. The loss function enforces cyclic consistency between robot video embedding are $\mathbf{z_{R}}=\{z_{R}^{1},z_{R}^{2},\dots,z_{R}^{T}\}$ and its demonstrator $\mathbf{z_{H}}=\{z_{H}^{1},z_{H}^{2},\dots,z_{H}^{T^{\prime}}\}$. Starting from a robot frame $z^{R}_{t}$, the TCC loss first computes $\mathbf{\alpha}$, a similarity distribution of $z^{R}_{t}$ with respect to the demonstrator’s embeddings, cycling to its soft-nearest neighbor $\tilde{z}_{H}=\sum_{t^{\prime}=1}^{T^{\prime}}\alpha_{t}z^{t^{\prime}}_{H}$. Then, $\tilde{z}_{H}$ cycles back to the robot video by again computing its similarity distribution $\mathbf{\beta}$ with respect to robot video embeddings to get its soft-nearest neighbor $\tilde{z}_{R}^{t}=\sum_{t=1}^{T}\beta^{t}z^{t}_{R}$. Formally, the cycling process is described as:
$\alpha_{k}=\frac{e^{-\|z_{R}^{t}-z_{H}^{t^{\prime}}\|^{2}}}{\sum_{t^{\prime}=1}^{T^{\prime}}e^{-\|z_{R}^{t}-z_{H}^{t^{\prime}}\|^{2}}}\quad\tilde{z}_{H}=\sum_{t^{\prime}=1}^{T^{\prime}}\alpha_{t}z^{t^{\prime}}_{H}\quad$
$\beta^{k}=\frac{e^{-\|\tilde{z}_{H}-z_{R}^{k}\|^{2}}}{\sum_{t=1}^{T}e^{-\|\tilde{z}_{H}-z_{R}^{t}\|^{2}}}\quad\tilde{z}_{R}^{t}=\sum_{t=1}^{T}\beta^{t}z^{t}_{R}$
The TCC loss function for a robot frame $z_{R}^{t}$ is the mean square error with its cycled-back frame $\tilde{z}_{R}^{t}$ as $l_{tcc}=||z_{R}^{t}-\tilde{z}_{R}^{t}||_{2}$. We define the video-level TCC distance function by summing over the frame-level losses $d(\mathbf{z_{R}},\mathbf{z_{H}})=\sum_{t=1}^{T}l_{tcc}(z_{R}^{t})$.
Using the optimal transport formulation, we measure distance by matching the distribution of two video embeddings. In contrast, the TCC distance attempts to establish strong one-to-one correspondences between the frames of two videos. We hypothesize that video retrieval using TCC distance can be inaccurate in two scenarios: 1) when the embeddings from robot and demonstrator tasks are not closely aligned, and 2) when multiple robot embeddings correspond to a single demonstrator embedding, such as when a robot arm performs a single task sequentially while the demonstrator executes two tasks simultaneously.
@@ V Experiments
Simulation Setup and Datasets. We test our approach on the Franka Kitchen simulator [41], in which a 7-DOF Franka arm can perform 7 different tasks. 580 robot trajectories, each executing 4 tasks sequentially, are used to train our models. Datasets of cross-embodiment videos are created in three different ways, progressively increasing the embodiment gap. Introduced by XSkill [5], The Sphere-Easy dataset replaces the robot model’s visual rendering in the simulator with a sphere that follows the gripper’s position, creating only a visual gap. To create mismatches in manipulation styles, we construct the Sphere-Medium dataset by handcrafting randomized motion primitives to complete tasks. For example, the robot drags the kettle object while the demonstrating agent lifts and carries it. Finally, the Sphere-Hard dataset further challenges the robot as the demonstrator performs two tasks simultaneously, mimicking how humans use two hands. Fig 2 (left) illustrates the simulator and the cross-embodiment datasets. Before training policies, all methods utilize a database of unpaired robot and demonstrator videos to pre-train visual representations. All policies are trained using Diffusion Policy [42]. More details are provided in the appendix.
Baselines. XSkill [5] simply conditions on robot videos during train-time, and uses its shared representation space to generalize zero-shot to inputs of demonstrator videos at test-time. OraclePairing [1, 2] is the gold-standard approach, assuming pairing between a demonstrator and each robot trajectory, enabling conditioning on the demonstrator at train time. Our approach, RHyME, finds a middle ground. Without pairing, it imagines demonstrator trajectories that perform the same tasks as a robot trajectory (Section IV-B). We compare two variants of our algorithm, RHyME-TCC and RHyME-OT, which differ in their distance functions used for retrieval. We also show how vision representations can be improved using short-horizon robot-demonstrator task pairs.
Evaluation and Success Metrics. At test-time, we provide the robot policy with demonstrator videos consisting of a composition of 4 tasks unseen at train-time. We measure Task Recall, which measures recall by counting the successfully completed tasks shown in the demonstrator’s video, and Task Imprecision, which measures imprecision and reports the percentage of tasks the robot attempts incorrectly—those not specified in the demonstrator’s video.
@@
Fig 2 reports task recall and imprecision rates across datasets. As the demonstrator’s actions visually and physically deviate further from those of the robot, policies trained with our framework RHyME consistently outperform XSkill. OraclePairing baseline consistently achieves greater than $90\%$ task recall without almost any task imprecision. The easiest dataset, Sphere-Easy, only exhibits a cosmetic gap between embodiments, keeping robot and demonstrator object movements the same. All methods have high task recall while maintaining precision. Still, we see that RHyME-OT has greater than $90\%$ task recall compared to XSkill’s $82\%$. This gap increases on the Sphere-Medium dataset ($72\%$ vs $42\%$) and becomes larger on the hardest dataset, Sphere-Hard.
Fig 3 (left) investigates this trend by probing the visual representations of the video encoder $\phi$, common across policies. We plot the image embeddings of the robot and demonstrator across three different tasks using a t-SNE plot. Sphere-Easy dataset, embeddings generated by both the robot and demonstrator executing the same tasks are clustered very closely. This explains why the XSkill policy can zero-shot to demonstration embeddings at test time. However, as soon as there are slight deviations between the robot and demonstrator, as seen in Sphere-Medium, its performance quickly deteriorates. Still, the embeddings of the same task are close enough for both RHyME algorithms to successfully retrieve the correct demonstration videos at train-time. The demonstrator performs two tasks simultaneously in Sphere-Hard. Consequently, demonstrator embeddings of two simultaneous tasks lie between the embeddings of the robot each of the tasks. RHyME-OT maintains its performance relative to baseline as it can successfully match sequences of robot embeddings with demonstration videos at train time. However, we observe that RHyME-TCC suffers a significant drop in task recall compared to RHyME-OT ($34\%$ vs $53\%$). More interestingly, this method has much higher task imprecision, even in the easier Sphere-Medium dataset. This suggests inaccurate retrieval of demonstration videos at train-time.
Figure 4: Demonstration Imagination We compute the Optimal Transport (OT) costs to measure the similarity between robot and demonstrator videos. We show the transport cost matrices, where brighter colors indicate costs. OT costs are lowest when tasks are the same between videos. Correct retrievals are highlighted in green.
@@
In the Sphere-Easy setting, where task embeddings are clustered well (Fig 3), the TCC distance is low between matching robot and demonstrator videos since embeddings cycle to and from other embeddings within the same cluster. In the harder settings where clusters are inherently ambiguous due to execution mismatches, mistakes in cycling-back are more frequent. Fig 3 (right) demonstrates a common failure mode of TCC in Sphere-Hard: both clips complete the same two tasks, but TCC yields high distance since the robot frame moving the kettle cycles to the demonstrator multi-tasking with the kettle and light switch, thus having a significant probability of incorrectly cycling back to the robot turning on the light switch. RHyME-TCC ultimately shows some improvements in task recall compared to XSkill on these dataset, but also has increased imprecision due to incorrect retrievals (Fig 2). Notably, RHyME-OT outperforms RHyME-TCC in both metrics across all datasets.
The key reason for this performance difference is that optimal transport computes distances by matching videos across a sequence of embeddings. Fig 4 visualizes the cost of the optimal transport plan between prompt robot clips and demonstration videos in the hard Sphere-Hard dataset. Comparing a robot clip doing two tasks (e.g. kettle and light), the transport cost across assignments is minimum only when compared to the demonstrator performing those same two tasks. TCC, on the other hand, attempts to establish one-to-one correspondences between the robot and demonstration frames, which are lacking in this dataset.
Figure 5: Performance Improves by Pairing Skills on Sphere-Hard. For both non-retrieval and retrieval-based methods, we see policies exhibit improved performance when supplied with vision representations fine-tuned with short-horizon robot-demonstrator snippet pairs and a contrastive optimal transport loss.
Q3. Does fine-tuning representations with short-horizon task pairs help?
In this experiment on the Sphere-Hard dataset, we use a small dataset of short-horizon task pairs and attempt to fine-tune the representation space to pull together sequences of embeddings ($\mathbf{z_{R}},\mathbf{z_{H}}$) that encode the same tasks. We employ the contrastive optimal transport-based task alignment loss outlined in Section IV-A, and find that fine-tuning with this objective improves vision representations across all baselines. We report results in Fig 5 for XSkill (fine-tuned) and RHyME-OT (fine-tuned). XSkill, which only trains its policy on robot embeddings, reaps more than a 25% task recall gain, up from 0%, when supplied with the fine-tuned representation space. However, RHyME (0% fine-tuned) still significantly outperforms all XSkill (fine-tuned) policies, as the embeddings seen across human and robot videos still differ in length and values. Ultimately, we find that comparing induced distributions over embeddings with optimal transport is a beneficial design choice for matching clips at a task-level in the face of execution mismatches. As the vision representations are fine-tuned on increasingly more task pairs, RHyME-OT’s task recall linearly increases. This is because task alignment helps retrieve video snippets more accurately for policy training.

### v2

@@ 5            $\pi_{\theta}\leftarrow\texttt{Update-Policy}(a_{t},\pi_{\theta}(s_{t},\mathbf{z_{R}}))$
Algorithm 2 RHyME: Train robot policy on both robot trajectories and imagined demonstrations
IV-A Training the Vision Encoder
We align the robot and demonstrator video embeddings in three ways: visually, temporally, and at the task level, all without requiring trajectory-level correspondences. We employ unsupervised losses $\mathcal{L}_{\rm vis}(\phi)$ and $\mathcal{L}_{\rm temp}(\phi)$ for visual and temporal alignment, following prior works [44, 45], and introduce an optional task alignment loss $\mathcal{L}_{\rm task}(\phi)$.
Figure 2: Cross-Embodiment Datasets (Left): We introduce three cross-embodiment demonstrators in the Franka Kitchen simulator with varying visual and execution mismatches. Sphere-Easy replaces the robot model with a sphere at the end effector location, inducing a visual gap. Sphere-Medium additionally deviates in manipulation style by moving objects along different trajectories. Sphere-Hard resembles humans, as the demonstrator is bimanual and parallelizes tasks. Results (Right): We report performance across all datasets on a range of baselines, showing that as the demonstrator’s execution visually and physically deviate further from those of the robot, policies trained with our framework RHyME consistently outperforms XSkill.
Visual Alignment. To align robot and demonstrator embeddings ($\mathbf{z_{R}}$, $\mathbf{z_{H}}$), we use SwAV [44], a self-supervised method that clusters images based on shared visual features. SwAV learns a set of $K$ prototype vectors, to which each image is assigned. The SwAV loss $\mathcal{L}_{\rm vis}(\phi)$ updates both the encoder and prototypes, aligning robot and demonstrator videos by clustering similar visual features.
Temporal Alignment. To align temporally adjacent frames in robot and demonstrator videos, we use Time Contrastive Loss [45]. This loss encourages embeddings of frames close in time to be similar. For each frame $z^{t}$, we define a positive set $\mathbf{z^{+}}$ of frames within a temporal window $w$, and a negative set $\mathbf{z^{-}}$ for frames outside this window. Using the contrastive loss $\mathcal{L}_{\rm temp}(\phi)$, we pull embeddings from the positive set closer and push negative set embeddings further apart, capturing temporal continuity across videos.
Task Alignment. Task-level alignment $\mathcal{L}_{\rm task}(\phi)$ is used when a small set of paired robot and demonstrator snippets is available. Unlike frame-level methods, this aligns video embeddings of the robot $\mathbf{z_{R}}$ and demonstrator $\mathbf{z_{H}}$. We compute the optimal transport distance $d(\mathbf{z_{R}},\mathbf{z_{H}})$ to measure the similarity between two sequences of video embeddings. We then apply a contrastive learning objective (INFO-NCE [46]) to pull matched embeddings closer and push different-task embeddings apart. The final task alignment loss is: $\mathcal{L}_{\rm task}(\phi)=-\sum\limits_{i}\frac{\text{exp}(-d(\mathbf{z_{R}}^{i},\mathbf{z_{H}}^{i}))}{\text{exp}(-d(\mathbf{z_{R}}^{i},\mathbf{z_{H}}^{i}))+\sum_{j\neq i}\text{exp}(-d(\mathbf{z_{R}}^{i},\mathbf{z_{H}}^{j}))}$
Our final loss function for training the visual encoder $\phi$ is:
$\mathcal{L}(\phi)=\lambda_{\rm vis}\mathcal{L}_{\rm vis}(\phi)+\lambda_{\rm temp}\mathcal{L}_{\rm temp}(\phi)+\lambda_{\rm task}\mathcal{L}_{\rm task}(\phi)$ (1)
where $\lambda_{\rm task}=0$ by default and non-zero only with access to short-horizon paired data (Fig. 5).
IV-B Training the Robot Policy
Training Overview. Algorithm 2 provides an overview of our approach to train a robot policy $\pi_{\theta}$ using both robot trajectories and imagined demonstration videos. The training process has two stages.
Stage 1: Create a Paired Dataset. For each robot trajectory $\mathbf{\xi_{R}}$ and video $\mathbf{v_{R}}$ in $D_{\rm robot}$, we encode the robot video into embeddings $\mathbf{z_{R}}=\phi(\mathbf{v_{R}})$ using the learned video encoder $\phi$. We then retrieve imagined demonstration embeddings $\mathbf{\hat{z}_{H}}$ by aligning $\mathbf{z_{R}}$ with demonstration snippets from the play dataset $D_{\rm play}$, through the function Imagine-Demo. This produces a paired dataset $D_{\rm paired}$ containing $(\mathbf{\hat{z}_{H}},\mathbf{z_{R}},\mathbf{\xi_{R}})$.
Stage 2: Train Policy on Paired Dataset. The policy $\pi_{\theta}$ is trained on the paired dataset $D_{\rm paired}$ in a hybrid fashion. For each element in $D_{\rm paired}$, we update the policy in two modes —Mode 1: The policy is conditioned on the robot video embeddings $\mathbf{z_{R}}$ to predict actions $\pi_{\theta}(a_{t}|s_{t},\mathbf{z_{R}})$, Mode 2: The policy is conditioned on the imagined demonstration embeddings $\mathbf{\hat{z}_{H}}$ to predict actions $\pi_{\theta}(a_{t}|s_{t},\mathbf{\hat{z}_{H}})$. By alternating between these two modes, the policy learns to generalize from both robot and imagined demonstration videos, enabling it to handle execution mismatches.
@@ IV-B Training the Robot Policy
Figure 3: Cross-Embodiment Vision Embeddings (Left): We use t-SNE to visualize cross-embodiment latent embeddings generated by robot and demonstrator when executing different tasks on all three datasets. TCC Failure Example (Right): The robot and video clip are equivalent, but specific frames have high TCC losses. For example, a frame showing the robot performing the ’kettle’ action has a high loss due to its nearest neighbor in the video performing both ’kettle’ and ’light’ actions. This frame cycles back to the robot performing ’light’, which is mismatched.
@@
Imagining Demonstration Videos. Algorithm 1 details the retrieval process for imaging a sequence of demonstration embeddings. We break the robot’s video into short-horizon windows and compare the embeddings with those from the play dataset, retrieving snippets with the lowest sequence-level distance. These retrieved snippets are concatenated to form an imagined long-horizon demonstration video. The key challenge is defining a distance function $d(\mathbf{z_{R}},\mathbf{z_{H}})$ that can handle video sequences of varying lengths. We propose two methods to compute this distance.
@@
Method 1: Optimal Transport Distance. We calculate the Wasserstein distance (Optimal Transport) [47] between the robot and demonstrator video embeddings, i.e., the cost of the optimal transport plan that transfers one sequence of video embeddings into another. The robot’s embedding distribution is defined as $\rho_{R}=\{1/T,1/T,\dots,1/T\}$, and the demonstrator’s embedding distribution is defined as $\rho_{H}=\{1/T^{\prime},1/T^{\prime},\dots,1/T^{\prime}\}$, where $T$ and $T^{\prime}$ are the lengths of the video sequences respectively. The cost function for the transport is $C\in\mathbb{R}^{T\times T^{\prime}}$ where $C^{ij}$ is the cosine distance between the robot embedding $z_{R}^{i}$ and the demonstration embedding $z_{H}^{j}$. Our goal is to find the optimal assignment $M\in\mathbb{R}^{T\times T^{\prime}}$ that transports the distribution from $\rho_{R}$ to $\rho_{H}$ while minimizing the cost of the plan. Formally, we need to find $M^{*}=\operatorname*{arg\,min}\limits_{M}\sum_{i}\sum_{j}C^{ij}M^{ij}$. After solving the optimal transport assignment, the distance function is the cost of the plan, i.e., $d(\mathbf{z_{R}},\mathbf{z_{H}})=\sum_{i}\sum_{j}C^{ij}M^{*ij}$. In practice, we optimize an entropy-regularized version of this problem to find an approximate solution efficiently using the Sinkhorn-Knopp algorithm [47].
@@
Method 2: Temporal-Cyclic Consistency (TCC) Distance. We calculate the TCC loss between human and robot videos following  [48] which computes cycle consistency between robot video embedding $\mathbf{z_{R}}=\{z_{R}^{1},z_{R}^{2},\dots,z_{R}^{T}\}$ and demonstrator $\mathbf{z_{H}}=\{z_{H}^{1},z_{H}^{2},\dots,z_{H}^{T^{\prime}}\}$. For each robot frame $z_{R}^{t}$, we first compute a similarity distribution $\mathbf{\alpha}$ of $z^{R}_{t}$ with respect to the demonstrator’s embeddings, to find a soft-nearest neighbor $\tilde{z}_{H}=\sum_{t^{\prime}=1}^{T^{\prime}}\alpha_{t}z^{t^{\prime}}_{H}$. Then, $\tilde{z}_{H}$ cycles back to the robot video by again computing its similarity distribution $\mathbf{\beta}$ with respect to robot video embeddings to get its soft-nearest neighbor $\tilde{z}_{R}^{t}=\sum_{t=1}^{T}\beta^{t}z^{t}_{R}$. The TCC distance for a robot frame $z_{R}^{t}$ is the mean square error with its cycled-back frame $\tilde{z}_{R}^{t}$ as $\ell_{\rm tcc}=||z_{R}^{t}-\tilde{z}_{R}^{t}||_{2}$. We define the video-level TCC distance function by summing over the frame-level losses $d(\mathbf{z_{R}},\mathbf{z_{H}})=\sum_{t=1}^{T}\ell_{\rm tcc}(z_{R}^{t})$.
We hypothesize that video retrieval using TCC distance can be inaccurate in two cases: (1) When robot and demonstrator embeddings differ due to variations in execution speed or style, leading to poor frame alignment. (2) When multiple robot embeddings correspond to a single demonstrator frame, as in sequential robot tasks versus parallel demonstrator actions, causing ambiguity in matching.
@@ V Experiments
Simulation Setup and Datasets. We evaluate our approach using the Franka Kitchen simulator [49], where a 7-DOF Franka arm performs $7$ different tasks. We generate $3$ cross-embodiment video datasets each progressively increasing the embodiment and execution mismatch, which contain $580$ long-horizon ($\sim$25 seconds) robot trajectories completing a sequence of 4 tasks and a bank of demonstrator play data ($>$ 3 hours) for training our models. First, in Sphere-Easy, we replace the robot’s visual rendering with a sphere following the gripper’s position, creating a visual gap between robot and demonstrator. Second, in Sphere-Medium, we introduce manipulation style mismatches by applying randomized motion primitives to the demonstrator, such as the robot dragging an object while the demonstrator lifts and carries it. Finally, in Sphere-Hard, we create a further divergence where the demonstrator perform two tasks simultaneously, similar to how humans use two hands. Fig. 2 (left) illustrates the cross-embodiment datasets. Before training, all methods pre-train visual representations on unpaired robot and demonstrator videos, and policies are trained using Diffusion Policy[50].
Baselines. XSkill [4] simply conditions on robot videos during train-time, and uses its shared representation space to generalize zero-shot to inputs of demonstrator videos at test-time. OraclePairing [7, 8] is the gold-standard approach, assuming an oracle pairs demonstration with a robot trajectory, enabling conditioning on the demonstrator at train time. Our approach, RHyME, finds a middle ground. Without pairing, it imagines demonstrator trajectories that perform the same tasks as a robot trajectory (Section IV-B) by exploiting sequence-level correspondences. We compare two variants of our algorithm, RHyME-TCC and RHyME-OT, which differ in their distance functions used for retrieval. We also show how vision representations can be improved using short-horizon robot-demonstrator task pairs.
Evaluation and Success Metrics. At test-time, we provide the robot policy with demonstrator videos consisting of composition of 4 tasks not seen during training. We evaluate one-shot imitation performance across 20 different demonstrator videos, rolling out the robot policy from the last 5 model checkpoints, yielding 100 total trials per dataset. We measure Task Recall, which measures recall by counting the successfully completed tasks shown in the demonstrator’s video, and Task Imprecision, which measures imprecision and reports the percentage of tasks the robot attempts incorrectly—those not specified in the demonstrator’s video.
@@
As the demonstrator’s actions visually and physically deviate further from those of the robot, policies trained with our framework RHyME consistently outperform XSkill (Fig. 2). The OraclePairing baseline consistently achieves greater than $90\%$ task recall with minimal task imprecision. In the prescence of merely a cosmetic gap in Sphere-Easy, all methods have high task recall while maintaining precision. Still, we see that RHyME-OT has higher $90\%$ task recall compared to XSkill’s $82\%$. This gap increases on the Sphere-Medium dataset ($72\%$ vs $42\%$) and becomes larger on Sphere-Hard ($53\%$ vs $1\%$).
Fig. 3 (left) investigates this trend by probing the visual representations of the video encoder $\phi$, common across policies. We plot the image embeddings of the robot and demonstrator across three different tasks using a t-SNE plot. In Sphere-Easy, robot and demonstrator embeddings are closely clustered by task execution, so the XSkill policy can zero-shot transfer to demonstration embeddings at test time, but quickly drops in performance as the execution mismatch increases. RHyME algorithms overcome this problem and successfully retrieve the correct demonstration videos at train-time by reasoning over sequences of embeddings. The demonstrator performs two tasks simultaneously in Sphere-Hard. Consequently, demonstrator embeddings during two simultaneous tasks lie between the individual task embedding clusters of the robot. RHyME-OT maintains its performance relative to baseline as a result, but we observe that RHyME-TCC suffers a significant drop in task recall compared to RHyME-OT ($34\%$ vs $53\%$). More interestingly, this method has much higher task imprecision, even in the easier Sphere-Medium dataset. This suggests inaccurate retrieval of demonstration videos at train-time.
Figure 4: Optimal Transport Distances. We measure the similarity between robot and demonstrator videos on the Sphere-Hard dataset by computing the cost of the Optimal Transport (OT) plans. The sum over the entire transport cost matrix costs yields the distance between videos. OT costs are lowest when tasks are the same between videos (highlighted by a tick mark).
@@
When task embeddings are well clustered (Sphere-Easy, Fig. 3), TCC correctly cycles back and forth within clusters to yield low distance for videos matching in task specification, and high distance otherwise. However, as task embedding clusters deviate due to execution mismatches, we observe inaccuracies: in Sphere-Hard (Fig. 3 (right)) when both clips complete the same two tasks, a bimanual task embedding lies in between two robot task clusters which, resulting in cycling-back to the incorrect robot frame. We see that RHyME-TCC has relatively high task recall with lower task imprecision (due to incorrect retrievals) relative to other methods. RHyME-OT performs strictly better across datasets (Fig. 2).
The key reason for this performance difference is that optimal transport computes distances by matching videos across a sequence of embeddings. Fig. 4 visualizes the cost of the optimal transport plan between prompt robot clips and demonstration videos in the hard Sphere-Hard dataset. Comparing a robot clip doing two tasks (e.g. kettle and light), the transport cost across assignments is minimum only when compared to the demonstrator performing those same two tasks. TCC, on the other hand, attempts to establish one-to-one correspondences between the robot and demonstration frames, which are lacking in this dataset.
Figure 5: Performance Improves by Pairing Skills on Sphere-Hard. For both non-retrieval and retrieval-based methods, performance improves when fine-tuned when the visual encoder is finetuned short-horizon robot-demonstrator snippet pairs using a contrastive optimal transport loss.
Q3. Does fine-tuning visual representations with task-equivalent pairs improve one-shot imitation?
Following Section IV-A, we perform a study assuming access to short-horizon task pairings across embodiments to fine-tune our representation space with the contrastive optimal transport-based task alignment loss $L_{task}(\phi)$ in the Sphere-Hard setting. We find that encouraging induced distributions over embeddings to be similar lifts the performance of both XSkill and RHyME-OT (Fig. 5), and scales up as more data is accessible for fine-tuning since task alignment helps accurately retrieve video snippets for policy training. Ultimately, comparing induced distributions over embeddings with optimal transport is a beneficial design choice for matching clips at a task-level in the face of execution mismatches, as RHyME-OT (0% fine-tuned) still significantly outperforms XSkill (fine-tuned).